Evening Shift July 24, 1999

B. Cox (cox@fnal.gov)
Sun, 25 Jul 1999 00:10:58 -0500 (CDT)

Shift Summary for Evening Shift, 26/Jul/99

On Shift: B. Cox, E. Cheu

Hours of beam available at all intensities: 0.
Hours of beam available at nominal intensity: 0.
Hours of E832 data (special runs not included): 0.
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time:

Tevatron cryo problems 15:30 - 21:40
Main Injector ground fault 21:41 - 24:00

========================================================================
Runs ended this shift:

Tapes taken to FCC this shift

None

========================================================================
Accesses during shift:

None

========================================================================
Overview of shift:

No beam

========================================================================
Description of Shift:

Run cosmics for lack of better things to do

22:15 Excitement - laser rms message breaks the monotony
however, apparently just a momentary glitch..
settle back in our chairs again.

23:00 Contribution to Kami find printer for <HP scope
hook up and with the combined
superior intelligences of the
crack shift personnel manage to
get it to work. Left in control
room waiting RK to feed the BA to
it.

Absolutely no beam...

Borrrringgg. This getting to be a habit

========================================================================
Follow up Notes

The X-terminals have problems due to a change of boot server
last week. Symptoms are that windows on the x-terminals apparently
freeze and/or are difficult to open. Apparently they cannot find the
fonts that they need. Problem will be dealt with on Monday. See
instructions in Log XX page 249 on how to recover.